PVS^{TM} License Agreement

Alicia Siciliano Computer Science Laboratory SRI International Menlo Park CA 94025 USA

                      Email: pvs-sri@csl.sri.com
                www: http://www.csl.sri.com/pvs.html

Phone: +1 (650) 859-3333 Fax: +1 (650) 859-2844

December 15, 2005

Please provide the information requested on pages 1, 4, and (if outside the USA) 5, sign, and return the entire agreement to the address or Fax number above.

There is no need to complete a new agreement if you already have a properly executed license for an earlier version of PVS.

Although not part of the agreement, we appreciate it if you will inform us at pvs-bugs@csl.sri.com of any bugs that you discover, and notify us at pvs-sri@csl.sri.com of interesting applications you have made of PVS and of any papers or reports describing your use of the system.

SRI INTERNATIONAL END-USE LICENSE AGREEMENT

SOFTWARE PROGRAM: PVS^{TM}

GRANT

SRI International (“SRI”) hereby grants to a royalty-free, non-exclusive, non-transferable license under the Licensed IP to use its ac-

enter business name

companying, above referenced software program (“Software”) and related documentation (“Documentation”) upon the terms and conditions set forth below:

Definitions

“Licensed Software” means the Software, and any larger applications in which such Software is incorporated.

“Licensed IP” means all patent claims, copyright, and trade secrets owned or freely licensable by SRI that are embodied in or necessarily practiced by the Software.

“Non-Commercial Purposes” means usage of Licensed Software (a) by an official governmental agency, strictly for non-commercial, public benefit purposes, or (b) solely for non-commercial research purposes in exchange for which no financially valuable consideration (including but not limited to sales or license revenue, service revenue, and advertising revenue, and whether in the form of monetary, equity, or other forms of consideration) is received, except for research funding received from an official government agency sponsor, a non-profit foundation, or similar research funding organization.

You may: (a) use or copy the Licensed Software solely for Non-Commercial Purposes;

(b) distribute copies of the Licensed Software solely for Non-Commercial Purposes and only if such copies are distributed free of charge to all recipients (including one copy free of charge to Licensor) and subject to this Agreement. Any copies of Licensed Software distributed under this Agreement must include a copy of the following notice to recipients prominently displayed:

This software is subject to the SRI International End-Use License Agreement for the PVS (tm) software package (the “License”). You may not use, copy, or distribute this software except in compliance with the License, and except after submitting a signed copy of the License to SRI. You may obtain a copy

     of the License at http://pvs.csl.sri.com/license.html or by sending email

to pvs-request@csl.sri.com. Software distributed under the License is distributed on an “AS IS” basis, WITHOUT WARRANTY OF ANY KIND, either express or implied. See the License for the specific language governing rights and limitations under the License. Portions of the software are Copyright (c) SRI International. All rights reserved. “PVS” is a trademark of SRI International, a California nonprofit public benefit corporation.

You may not: modify the Software, or decompile, disassemble, or otherwise copy the Software in modified form; or publish, distribute, or transfer copies of the Licensed Software to any third party except under the limited terms set forth above. These examples are by way of illustration, not limitation; except as expressly set forth above (under “You may”), no other license rights are implied or otherwise granted under this License Agreement.

TITLE

SRI retains title, ownership rights, and intellectual property rights in and to the Software and Documentation. The Software is protected by the copyright laws of the United States and international copyright treaties.

DISCLAIMER OF WARRANTY

Since this Software is a research tool still in the development stage and is provided free of charge, the Software and Documentation are provided on an “AS IS” basis, WITH-OUT WARRANTY OF ANY KIND, INCLUDING WITHOUT LIMITATION THE WAR-RANTIES OF MERCHANTABILITY OR FITNESS FOR ANY PARTICULAR PUR-POSE OR THAT THE USE OF THE SOFTWARE OR DOCUMENTATION WILL NOT INFRINGE ANY PATENTS, COPYRIGHTS, TRADEMARKS, OR OTHER RIGHTS. The entire risk as to the quality and performance of the Software and/or Documentation is borne by you. Should the Software and/or Documentation prove defective, you and not SRI assume the entire cost of any services and repair. This disclaimer of warranty constitutes an essential part of the agreement.

LIMITATION OF LIABILITY

UNDER NO CIRCUMSTANCES AND UNDER NO LEGAL THEORY, TORT, CON-TRACT, OR OTHERWISE, SHALL SRI, SRI’S AGENTS, OFFICERS, ASSISTANTS AND EMPLOYEES EITHER IN THEIR INDIVIDUAL CAPACITIES OR BY REASON OF THEIR RELATIONSHIP TO SRI AND SRI’S SUCCESSORS, BE HELD LIABLE TO YOU OR ANY OTHER PERSON FOR ANY INDIRECT, SPECIAL, INCIDENTAL, OR CONSEQUENTIAL DAMAGES OF ANY CHARACTER INCLUDING, WITHOUT LIMITATION, DAMAGES FOR LOSS OF GOODWILL, WORK STOPPAGE, COM-PUTER FAILURE OR MALFUNCTION, OR ANY AND ALL OTHER COMMERCIAL DAMAGES OR LOSSES. IN NO EVENT SHALL SRI BE LIABLE FOR COSTS OF PRO-CUREMENT OF SUBSTITUTE OR REPLACEMENT PRODUCTS OR SERVICES.

NO SUPPORT, NO UPDATES

This license does not grant you any right to any Software support, enhancements or updates.

TERMINATION

This license will terminate automatically if you fail to comply with the limitations described above. On termination, you must destroy all copies of the Software.

EXPORT RESTRICTIONS

Each party hereby acknowledges that the rights and obligations of this Agreement are subject to the laws and regulations of the United States relating to the export of products and technical information. Without limitation, each party shall comply with all such applicable laws and regulations.

U.S. GOVERNMENT RESTRICTED RIGHTS

If the Software or any accompanying documentation is used or acquired by or on behalf of any unit, division or agency of the United States Government, this provision applies. The Software and any accompanying documentation is provided with RESTRICTED RIGHTS. The use, modification, reproduction, release, display, duplication or disclosure thereof by or on behalf of any unit, division or agency of the Government is subject to the restrictions set forth in subdivision (c)(1) of the Commercial Computer Software-Restricted Rights clause at 48 CFR 52.227-19 and the restrictions set forth in the Rights in Technical Data-Non-commercial Items clause set forth in 48 CFR 252.227-7013. The contractor/manufacturer of the Software and accompanying documentation is SRI International, 333 Ravenswood Avenue, Menlo Park, California 94025, USA.

MISCELLANEOUS

This agreement represents the complete agreement concerning this license between the parties and supersedes all prior agreement and representations between them. It may be amended only by a writing executed by both parties. If any provision of this Agreement is held to be unenforceable for any reason, such provision shall be reformed only to the extent necessary to make it enforceable. This agreement shall be governed by and construed in accordance with the laws of the State of California, without regard to the conflicts of law principles thereof, and shall not be governed by the United Nations convention on contracts for the international sale of goods.

SIGNATURES

The licensed material is provided for the use of LICENSEE at the following

SITE/FACILITY only:  Enter location where PVS will be used

Facility: Address:

Enter name and address of person for correspondence

Name: Address:

(leave address blank if same as above)

Phone: Fax: Email:

In the spaces below on the right, enter the name of your business or institution. The person signing below the business name should be someone who has the legal authority to sign for your business or institution. That person’s position of authority (title) in the business or institution should be entered on the next line. Return the completed license to SRI at the address on the cover; SRI will countersign on the left below and return to you at the address you enter above.

LICENSOR: SRI International

By

Title

Date LICENSEE:

By

Title

Date END-USE STATEMENT

enter business name enter signature enter business title

Use of PVS outside the USA may be subject to US Government export licensing requirements. The requirements are dependent on the use to be made of PVS. Overseas users must therefore submit with their completed license a statement of the end-use to which they plan to put the software.

If it applies to your use of PVS, we suggest you use the following form of words:

PVS will be used to conduct basic and applied research. It is intended that the results of the research will be made generally available, and published should they merit it.

If your use of PVS will concern something other than research, or if the results will be restricted for proprietary or national security reasons, then please provide explicit details in your end-use statement.

Please write or type your end-use statement below.